OpenStudy (anonymous):

How do you get the RC time constant of a charging circuit given by voltages values and time?

OpenStudy (jamesj):

Well, if it is a simple RC circuit, then you know that the voltage over the capacitor is of the form V(t) = V(1-e^(-t/RC)), where V is the voltage difference over the voltage source. Thus -t/RC =ln ( 1 - V(t)/V ) Given values of the V(t) at a set of values for t, you can run a simple regression to estimate the value of 1/RC.